please help i dont understand how to do this
•Model A takes 2 hours to assemble.
•Model B takes 3 hours to assemble.
•Model A costs $25 to make per bike.
•Model B costs $30 to make per bike.
If your company has a total of 34 hours and $350 available per day for these two models, how many of each model can be made in a day?

Answer by mananth(16949)

Model A = x numbers
Model B = y numbers
cost equation
25x+30y=350.............1
time equation
2x+3y=34............2
multiply equation 2 by -10 and add to equation 1
-20x-30y+25x+30y=-340+350
5x=10
x= 2 numbers Model A
..
2*2+3y=34
4+3y=34
3y=34-4
3y=30
y=10 numbers Model B